OpenStudy (anonymous):

v^2-4=0 solve and check

OpenStudy (jamesj):

So v^2 = 4. hence what values can v have?

OpenStudy (anonymous):

v^2 - 4 = 0 (v+2)(v-2) = 0 v = 2 or - 2